# Mechanisms vs. Outcomes: Probing for Syntax Fails to Explain Performance on Targeted Syntactic Evaluations

### v1

@@ 7 Limitations
Our evaluation only uses English grammatical benchmarks and could usefully be extended to other languages, but a failure of probe predictive power on English suggests the same result would hold for other languages. Moreover, while we design our syntax probes to be lightweight and to learn to extract syntax with different training objectives, they do not encompass all possible syntactic probing tasks or architectural variants.
Recent probing work aims to make causal claims between probed representations and observed model downstream behavior Belinkov (2022). By intervening to modify these representations, such work investigates if these representations are actually “used” by the model for specific tasks. Our work does not perform causal interventions; instead our focus is the viability of the mechanism conveyed by probing – syntactic knowledge can be linearly decoded from contextual representations – at several experimental granularities. Since our results are primarily negative, there is no discovered correlation to investigate deeper for causality anyway.
@@ Appendix A Models
Pythia 410M Biderman et al. (2023) EleutherAI/pythia-410m
@@
GPT-J 6B Wang and Komatsuzaki (2021) EleutherAI/gpt-j-6b
Falcon 7B Almazrouei et al. (2023) tiiuae/falcon-7b
Llama 3.2 1B Grattafiori et al. (2024) meta-llama/Llama-3.2-1B
@@
Llama 2 7B Touvron et al. (2023) meta-llama/Llama-2-7b-hf
@@
Mistral 7B Jiang et al. (2023) mistralai/Mistral-7B-v0.1
Gemma 2B Team et al. (2024a) google/gemma-2b
Gemma 2 2B Team et al. (2024b) google/gemma-2-2b
OLMo 1B Groeneveld et al. (2024) allenai/OLMo-1B-hf
OLMo 7B Groeneveld et al. (2024) allenai/OLMo-1.7-7B-hf
Phi 2 2.7B Gunasekar et al. (2023) microsoft/phi-2
@@
ELECTRA-Small 14M Clark et al. (2020) google/electra-small-discriminator
@@
RoBERTa 125M Liu et al. (2019) FacebookAI/roberta-base
@@
DeBERTa v2 900M He et al. (2021) microsoft/deberta-v2-xlarge
@@ Appendix D Critical Tree Edges for BLiMP Phenomena
We investigate if the syntax probe’s accuracy in predicting a critical edge we define on a per-paradigm basis in the acceptable sentence’s parse tree is consistent with the model’s outcome of assigning the acceptable sentence a higher log probability than the unacceptable one. The critical edge involves the minimal difference in the pair’s sentences. In the following sections, we justify the critical edges for the subject-verb agreement and filler-gap paradigms that we experiment with:
  • Subject-verb agreement
@@
D.1 Subject-Verb Agreement
@@
Agreement on the verb ‘aggravate’ is determined by its subject, the plural noun ‘prints’. This is represented in parse trees as an edge connecting ‘prints’ and ‘aggravate,’ labeled nsubj. Fig 8 shows the parse tree for the acceptable sentence. If a model’s syntactic representations are related to its success on subject-verb agreement, nsubj is the critical dependency to evaluate – a model successfully representing this edge means that it is minimally aware of what noun is its subject, and thus, which noun it should use to determine whether ‘aggravate’ or ‘aggravates’ should be assigned a higher log probability (see Clark et al., 2019; Jian and Reddy, 2023 for similar discussions of syntactic attention heads).
Under the UD formalism whose parses Stanza produces, the verb which agrees with the subject is not always the head of the nsubj relation in a sentence. For example, in Fig 10 the verb which agrees with the subject is the auxiliary ‘have’, but the UD parse assigns ‘alarmed,’ which does not agree, as the head. For these critical edge experiments only, we filter out all incongruent cases where UD assigns no edge between the subject and critical verb, which leaves 236 valid minimal pairs for distractor agreement relational noun and 345 for distractor agreement relative clause.
{dependency}
[edge style=blue!60!black,very thick] {deptext}[column sep=.5cm, row sep=.1ex] DT & NNS & IN & DT & NN & VBP & NNP & .
The & prints & of & every & vase & aggravate & Nina & .
[…]
@@
{dependency}
[edge style=blue!60!black,very thick] {deptext}[column sep=.5cm, row sep=.1ex] DT & NNS & WDT & VBD & VBG & DT & NN & VB & RB & .
All & libraries & that & were & aggravating & a & legislature & joke & around & .
[…]
@@
{dependency}
[edge style=blue!60!black,very thick] {deptext}[column sep=.5cm, row sep=.1ex] DT & NNS & IN & NN & VBP & VBN & NNP & .
The & plays & about & art & have & alarmed & Mitchell & .
[…]
@@
{dependency}
[edge style=blue!60!black,very thick] {deptext}[column sep=.5cm, row sep=.1ex] NNP & VBD & VBN & WP & DT & NN & VBD & .
Marcus & had & remembered & who & some & lady & disliked & .
[…]
@@
{dependency}
[edge style=blue!60!black,very thick] {deptext}[column sep=.2cm, row sep=.1ex] NNP & VBD & WP & DT & NNS & WDT & VBD & JJ & MD & VB & RP & .
Brenda & knew & what & all & bananas & that & went & bad & can & look & like & .
[…]
@@ D.2 Filler-gap
We establish the critical edge for wh. vs. that with gap and wh. vs. that with gap long distance in the same way. Fig 11 and Fig 12 show sample acceptable parse trees for wh. vs. that with gap and wh. vs. that with gap long distance. The critical edge is the edge between the wh-word and the verb it is the direct (obj) or oblique object (obl) of; any other case is filtered out. After filtering, there are 972 valid minimal pairs for wh. vs. that with gap and 885 for wh. vs. that with gap long distance.
@@ Appendix E Probe Training and Comparison
Figure 13: Penn Treebank test set head word probe UAS for each layer of a sample of our models. We train a probe on each layer of the model using the train and validation splits and select the probe for the layer with the best test set accuracy – indicated in the plot with a star icon – for BLiMP evaluation. For most models, the best layer is in the first half of the model.
Figure 14: Correlating $g^{\text{head}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ BLiMP attachment scores. Across all models, $g^{\text{head}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ syntax probes show a significant, moderately positive correlation in accuracy averaged across BLiMP test suites. Despite having an additional directional aspect, $g^{\text{head}}_{\phi}$ trains similarly to $g^{\text{struct}}_{\phi}$.
Figure 15: Correlating $g^{\text{ortho}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ BLiMP attachment scores. Across all models, $g^{\text{ortho}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ syntax probes show a strongly positive correlation in accuracy averaged across BLiMP test suites. Despite the larger capacity of the orthogonal probe, performance largely mirrors its simpler variant, suggesting that the extra dimensions are not critical to learning the parse tree probing task.
@@ Appendix F Full BLiMP and per-Phenomenon OLS Regressions
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
@@
Table 5: Comparison of simple and multiple linear regressions with $x_{1}=$ $g^{\text{struct}}_{\phi}$ UUAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$, extending Table 1. Per-phenomenon rows are sorted in order of increasing simple regression corrected $p$-value. Anaphor agreement is the only one with statistical significance even after incorporating the control. The other phenomena show substantial variation in the strength and direction of association. In particular, we note that the adjusted $R^{2}$ jumps for s-selection and irregular forms when adding the control. The significant $p$-value of the irregular forms LRT is concurrent with Fig 5 showing the control probe to be a moderate predictor of irregular forms minimal pairs.
@@
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
@@
Table 6: Comparison of simple and multiple linear regressions with $x_{1}=$ $g^{\text{ortho}}_{\phi}$ UUAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$, extending Table 2. Per-phenomenon rows are sorted in order of increasing simple regression corrected $p$-value.
@@
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
@@
Table 7: Comparison of simple and multiple linear regressions with $x_{1}=$ $g^{\text{head}}_{\phi}$ UAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$, extending Table 3. Per-phenomenon rows are sorted in order of increasing simple regression corrected $p$-value. Having 12/13 phenomena with $p$-value 1.0 is a striking result; moreover, comparing with $g^{\text{struct}}_{\phi}$ regressions in Table 5, $g^{\text{head}}_{\phi}$ shows relatively much weaker ability to fit anything at all.
@@
Table 8: Simple linear regression results for the structural probe control (trained on the best layers for $g^{\text{struct}}_{\phi}$) $x_{1}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$. Per-phenomenon rows are sorted in order of increasing corrected $p$-value. Irregular forms has statistical significance and substantially higher $R^{2}$ than most of the syntax probe simple regressions for any phenomena (Table 5 and Table 7). This critical result shows that even a non-syntactic control can achieve a significant positive association with one syntactic benchmark, which thus urges caution when interpreting the statistical signifiance of anaphor agreement for $g^{\text{struct}}_{\phi}$.
@@
Appendix G Subject-Verb Agreement & Filler-Gap Experiments

### v2

@@ 7 Limitations
Our evaluation only uses English grammatical benchmarks and could usefully be extended to other languages. Extending our work to multilingual benchmarks would strengthen the study. This would require formalization of a unified evaluation framework to handle differing linguistic phenomena across languages. Moreover, while we design our syntax probes to be lightweight and to learn to extract syntax with different training objectives, they do not encompass all possible syntactic probing tasks or architectural variants. Other probe architectures could yield different results. Due to compute limitations, we also do not train probes on models larger than 8B parameters.
@@ Appendix A Models
Pythia 410M biderman2023pythia EleutherAI/pythia-410m
@@
GPT-J 6B gpt-j EleutherAI/gpt-j-6b
Falcon 7B almazrouei2023falconseriesopenlanguage tiiuae/falcon-7b
Llama 3.2 1B grattafiori2024llama3herdmodels meta-llama/Llama-3.2-1B
@@
Llama 2 7B touvron2023llama2openfoundation meta-llama/Llama-2-7b-hf
@@
Mistral 7B jiang2023mistral7b mistralai/Mistral-7B-v0.1
Gemma 2B gemmateam2024gemmaopenmodelsbased google/gemma-2b
Gemma 2 2B gemmateam2024gemma2improvingopen google/gemma-2-2b
OLMo 1B Groeneveld2023OLMo allenai/OLMo-1B-hf
OLMo 7B Groeneveld2023OLMo allenai/OLMo-1.7-7B-hf
Phi 2 2.7B gunasekar2023textbooksneed microsoft/phi-2
@@
ELECTRA-Small 14M clark2020electrapretrainingtextencoders google/electra-small-discriminator
@@
RoBERTa 125M DBLP:journals/corr/abs-1907-11692 FacebookAI/roberta-base
@@
DeBERTa v2 900M he2021deberta microsoft/deberta-v2-xlarge
@@ Appendix D Critical Tree Edges for BLiMP Phenomena
We investigate if the syntax probe’s accuracy in predicting a critical edge we define on a per-paradigm basis in the acceptable sentence’s dependency parse tree is consistent with the model’s outcome of assigning the acceptable sentence a higher log probability than the unacceptable one. The critical edge involves the minimal difference in the pair’s sentences. In the following sections, we justify the critical edges for the subject–verb agreement and filler–gap paradigms that we experiment with:
  • Subject–verb agreement
@@
D.1 Subject–Verb Agreement
@@
Agreement on the verb ‘aggravate’ is determined by its subject, the plural noun ‘prints’. This is represented in parse trees as an edge connecting ‘prints’ and ‘aggravate,’ labeled nsubj. Fig D.1 shows the parse tree for the acceptable sentence. If a model’s syntactic representations are related to its success on subject–verb agreement, nsubj is the critical dependency to evaluate – a model successfully representing this edge means that it is minimally aware of what noun is its subject, and thus, which noun it should use to determine whether ‘aggravate’ or ‘aggravates’ should be assigned a higher log probability (see clark2019bert; jian-reddy-2023-syntactic for similar discussions of syntactic attention heads).
Under the UD formalism whose parses Stanza produces, the verb which agrees with the subject is not always the head of the nsubj relation in a sentence. For example, in Fig D.1 the verb which agrees with the subject is the auxiliary ‘have’, but the UD parse assigns ‘alarmed,’ which does not agree, as the head. For these critical edge experiments only, we filter out all incongruent cases where UD assigns no edge between the subject and critical verb, which leaves 236 valid minimal pairs for distractor agreement relational noun and 345 for distractor agreement relative clause.
@@ D.2 Filler-gap
We establish the critical edge for wh. vs. that with gap and wh. vs. that with gap long distance in the same way. Fig D.1 and Fig D.1 show sample acceptable parse trees for wh. vs. that with gap and wh. vs. that with gap long distance. The critical edge is the edge between the wh-word and the verb it is the direct (obj) or oblique object (obl) of; any other case is filtered out. After filtering, there are 972 valid minimal pairs for wh. vs. that with gap and 885 for wh. vs. that with gap long distance.
@@ Appendix E Probe Training and Comparison
Figure 13: Penn Treebank test set $g^{\text{head}}_{\phi}$ UAS for each layer of a sample of our models. We train a probe on each layer of the model using the train and validation splits and select the probe for the layer with the best test set accuracy – indicated in the plot with a star icon – for BLiMP evaluation. For most models, this occurs in the first half.
Figure 14: Correlating $g^{\text{head}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ BLiMP attachment scores. Across all models, $g^{\text{head}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ syntax probes show a significant, moderately positive correlation in accuracy averaged across BLiMP paradigms. Despite having an additional directional aspect, $g^{\text{head}}_{\phi}$ trains similarly to $g^{\text{struct}}_{\phi}$.
Figure 15: Correlating $g^{\text{ortho}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ BLiMP attachment scores. Across all models, $g^{\text{ortho}}_{\phi}$ and $g^{\text{struct}}_{\phi}$ syntax probes show a strongly positive correlation in accuracy averaged across BLiMP paradigms. Despite the larger capacity of the orthogonal probe, performance largely mirrors its simpler variant, suggesting that the extra dimensions are not critical to learning the dependency tree probing task.
@@ Appendix F Full BLiMP and per-Phenomenon OLS Regressions
[…]
@@
(a) $x_{1}=$ $g^{\text{struct}}_{\phi}$ UUAS.
@@
[…]
@@
(b) $x_{1}=$ $g^{\text{ortho}}_{\phi}$ UUAS.
@@
[…]
@@
(c) $x_{1}=$ $g^{\text{head}}_{\phi}$ UAS. That 12 out of 13 phenomena have pp-value 1.0 is a striking result.
Table 3: Comparison of simple and multiple regression statistics for all BLiMP linguistic phenomena. Per-phenomenon rows are sorted in order of increasing simple regression corrected $p$-value. Phenomena show substantial variation in strength of association. Zeroing into Subtable 2(a), we note the adjusted $R^{2}$ jumps for s-selection and irregular forms after adding the control. The significant $p$-value of the irregular forms LRT concurs with Fig 5.
@@
Table 4: Simple linear regression results for the structural probe control (trained on the best layers for $g^{\text{struct}}_{\phi}$) $x_{1}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$. Per-phenomenon rows are sorted in order of increasing corrected $p$-value. Irregular forms has statistical significance and substantially higher $R^{2}$ than most of the syntax probe simple regressions for any phenomena (Table 2(a) and Table 2(c)). This critical result shows that even a non-syntactic control can achieve a significant positive association with one syntactic benchmark, which thus urges caution when interpreting the statistical signifiance of anaphor agreement for $g^{\text{struct}}_{\phi}$.
@@
Appendix G Subject–Verb Agreement & Filler–Gap Experiments